OpenStudy (anonymous):

Help PLEASE ! : - Evaluate f(x)=2xplus4 for the following .

OpenStudy (anonymous):

what is the following?

OpenStudy (anonymous):

f(1)

OpenStudy (anonymous):

f(4)

OpenStudy (anonymous):

just substitute

OpenStudy (anonymous):

and f(-2)

OpenStudy (anonymous):

i d k how to do that , i hate math .

OpenStudy (anonymous):

f(1)=2*(1)+4

OpenStudy (anonymous):

and so on

OpenStudy (anonymous):

do the same for f(4) f(-2)

OpenStudy (anonymous):

lolthos is hard

OpenStudy (anonymous):

just do like i just say i will help u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me know if u need other help hv a nive day:)

OpenStudy (anonymous):

ok ( :

OpenStudy (anonymous):

Evaluatef(x)=1/4x-4 ?

OpenStudy (anonymous):

what r the values of x

OpenStudy (anonymous):

f(4) and f(-8)

OpenStudy (anonymous):

ok f(4)=1/4(4)-4 and then just calculate ok the same for f(-8)

OpenStudy (anonymous):

ok ok thankkksss !!(:

OpenStudy (anonymous):

no problem hv a nice day:)

OpenStudy (anonymous):

you too !!!